Shift report 20000725 16:00-24:00

From: http account (http@pii3.brahms.bnl.gov)
Date: Tue Jul 25 2000 - 23:59:33 EDT

  • Next message: http account: "Shift report 20000726 00:00-08:00"

    ------------------------------------------------------------
    Date: Tue Jul 25 23:59:32 2000, Shift: 16:00-24:00
    Supervisor: kh, with: yb
    Log book#3 pages 78 - 80
    
    Shift Summary:
    ------------------------------------------------------------
    No beam during entire shift.
    
    There was a quench interlock during a large part of the
    shift.
    
    Discovered a problem with generated ntuples in that
    the ZDC high and low gain ADC's gave exactly the same
    spectra.  Fixed this problem with BrNtuple by relinking
    the program.
    
    Unfortunately this fix did not make the single neutron
    peaks reappear.
    
    TPC voltages off during entire time.
    ------------------------------------------------------------
    



    This archive was generated by hypermail 2b29 : Wed Jul 26 2000 - 00:03:03 EDT